Consider a project that costs $250 now and is expected to generate $98 in net revenues at the end of each of the next three years. If the MARR is 5%, the future worth of this project is __________.
Answer #1

Answer

Future worth =FV of benefits -FV of costs

FV=PV*(1+i)^n or A*(F|A,i,n)

Future value =98*(F|A,5%,3)-250*(1.05^3)

=98*3.152-289.40625

=19.48975

The future value of the project is $19.49